Ammonium nitrate is produced when ammonia is combined with concentrated nitric acid as shown in the following balanced equation: NH,(aq) + HNO,(aq) — NH,NO)(aq) In an experiment, 404.2 g of ammonium nitrate is produced when 112.5 g of ammonia and excess nitric acid are combined. Determine the percentage yield of the product.

1
Expert's answer
2022-03-27T10:32:16-0400

NH3(aq) + HNO3(aq) — NH4NO3(aq)

nt (NH4NO3) = nt (NH3) = 112.5 g/ 17 g/mol = 6.62 mol

mt (NH4NO3) = nt (NH4NO3)* M(NH4NO3) = 6.62 mol* 80 g/mol = 529.6 g

mpr (NH4NO3) = 404.2 g

w(NH4NO3) = mpr (NH4NO3)/ mt (NH4NO3) = 404.2 g/ 529.6 g = 0.76 (76%)

Answer: 0.76 (76%)